OpenStudy (anonymous):

two stores had a combined total of 210 teddy bears. after the first store sold 1/2 of its teddy bears and the second store sold 1/3 of its teddy bears, each store had the same number of teddy bears left. how many bears were sold from the two stores combined??????

OpenStudy (shane_b):

x+y=210 (1/2)x=(2/3)y Now just solve the system of equations for x and y.

OpenStudy (anonymous):

i tried that and i could not get it to work out right how do you do it?

OpenStudy (shane_b):

Here are the two equations: \[x+y=210\]\[(\frac{1}{2})x=(\frac{2}{3})y\]I'll solve it by substituting the second equation into the first. Solving the second equation for x: \[(\frac{1}{2})x=(\frac{2}{3})y\]Divide both sides by 1/2 to get:\[x=\frac{4}{3}y\] Now, plug that back in to the first equation to get:\[\frac{4}{3}y+y=210\]\[\frac{7}{3}y=210\]\[y=90\] Now that you know y = 90 and x+y=210, x must be 120. If anything is still unclear please let me know :)